OpenStudy (anonymous):

timmy hits a ball at an angle 14.8 degrees above the horizontal . his friend lou returns the ball at the same level from which timmy originally hit. the time hang time of the ball is measured to be 1.5 seconds. what was the initial speed of the ball when timmy hit it? WHat was the horizontal displacement of the ball travelled when lou returned it?

OpenStudy (turingtest):

The only force acting on the ball is gravity. This force is only in the y-direction. The kinematic equations for the y-direction are then\[a_y=-g\]\[v_y=v_{iy}-gt\]\[s_y=v_{iy}t-\frac12gt^2\]There is no force in the x-direction, so there is no acceleration. The kinematics in the x-direction are\[a_x=0\]\[v_x=v_{ix}\]\[s_x=v_{ix}t\]You are given the time, so you can find the horizontal displacement quite easily if you know vx. The x and y component of the velocity is given by\[v_x=|v|\cos\theta\]\[v_y=|v|\sin\theta\]where |v| is the magnitude of the velocity. Play around with these equations and try to get enough pieces together to find |v|. Once you have that you can get vx, which means you can get horizontal displacement. That's the best way to practice these problems.